# On Achievable Spectral Efficiency Using Adaptive Transmission Over Terrestrial Coherent FSO Links

### v9

@@ II-B Noise Sources
Two fundamental sources of noise interfere in the process of photodetection in FSO receivers: i) the shot noise due to photodetection itself [41], and ii) the thermal noise originating in the receiver circuitry. The shot noise is due to the information optical radiation and the ambient light incident on the receiver lens, and its power grows linearly with the total incident optical power. Importantly, both ambient shot noise and thermal noise are signal-independent, as they remain statistically uncorrelated with the information-carrying optical signal [42].
The IM-DD photodetector output consists of the desirable photocurrent component mapped one-to-one to the intensity of the incoming (information) optical signal, signal-dependent shot noise, ambient shot noise, and thermal noise currents. The operating mode of an IM-DD receiver is determined by the relative strengths of these noise components.
In a coherent HD receiver, in contrast, an optical local oscillator (OLO) signal is mixed with the received signal prior to photodetection. This introduces an OLO-dependent shot noise current in addition to signal-dependent shot noise, ambient shot noise, and thermal noise currents. By ensuring a large OLO optical power, the OLO shot noise component can be made dominant compared to other noise components; this is in fact considered to be the de facto mode of operation of an (ideal) coherent HD receiver [10][42][43].
@@
Remark on the Statistical Characterization of Shot Noise: When the number of received photons per symbol interval is low-as is typical in inter-satellite and ground-to-satellite optical links-the shot noise is accurately modeled using a Poisson distribution [44]. In contrast, for high photon-count regimes, as encountered in most terrestrial FSO applications, a Gaussian approximation is sufficiently accurate. For a rigorous justification of the Gaussian model for high-intensity shot noise, refer [44][45].
II-C Coherent HD Receiver Model
@@
Let $x_{\mathrm{\scriptscriptstyle HD}}\in\mathbb{C}$ denote the transmitted symbol for the coherent HD system respectively. We remind the reader that the baseband (information) signal in a coherent HD scheme represents a complex optical field, whereas the baseband signal in an IM-DD scheme represents optical intensity. The transmitted beam power is constrained by a fixed average power constraint denoted by $\mathbb{E}\left[|x_{\mathrm{\scriptscriptstyle HD}}|^{2}\right]\leqslant P_{\mathrm{avg}}$. As justified in the introduction, owing to millisecond-scale coherence times and hundreds of gigahertz coherence bandwidths, a frequency-flat block fading model is well-justified for a terrestrial optical fading channel subjected to time-varying irradiance fluctuations: each fading state is assumed to span an identical duration, corresponding to a fixed coherence time, with i.i.d. fading realizations across blocks. The block-wise channel state is considered to be perfectly known at both the transmitter and the receiver at all times. In practice, this state needs to be estimated at the receiver only once at the beginning of each block, and conveyed to the transmitter over an error-free feedback link. Within this framework, we consider a FSO transmitter that performs signal adaptation on a block-by-block basis. Adhering to this understanding, the received signal model for the HD detection-based FSO receiver system is developed next.
The complex baseband symbol in a coherent HD receiver is described by
$y_{\mathrm{\scriptscriptstyle HD}}\,=\,h\,x_{\mathrm{\scriptscriptstyle HD}}\,+\,\,w_{\mathrm{\scriptscriptstyle OLO}}\,+\,w_{\mathrm{\scriptscriptstyle OTH}}$ (10)
@@
where $y_{\mathrm{\scriptscriptstyle HD}}\ \in\ \mathbb{C}$ represents its output, $h\in\mathbb{C}$ the fluctuation in the received optical field, $w_{\mathrm{\scriptscriptstyle OLO}}\sim\mathcal{CN}(0,\sigma^{2}_{\mathrm{\scriptscriptstyle OLO}})$ the OLO shot noise, and $w_{\mathrm{\scriptscriptstyle OTH}}$ the total AWGN from all the sources other than the OLO [13, Ch. 3]. As noted earlier, the OLO shot noise is the only dominant noise of consequence. Hence, conditioned on $h$ or more appropriately on $I:=|h|^{2}$, the received SNR is
$\gamma_{\mathrm{\scriptscriptstyle HD}}\,=\,{I\,|x_{\mathrm{\scriptscriptstyle HD}}|^{2}}/{\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}}.$ (11)
@@
III Achievable Spectral Efficiency With Adaptive Optical Beam Transmission
In an IM-DD FSO system, we notice that the transmitted signal represents optical intensity, and thus must be non-negative. On the other hand, the $x_{\mathrm{\scriptscriptstyle HD}}$ scheme has no such constraint and can utilize two degrees of freedom. As a result, the classical Shannon limit on spectral efficiency for the AWGN channel applies to the HD scheme, but not to the IM-DD system (see [46]). Conditioned on the received SNR in (11), the achievable SE of the coherent optical channel (in nats/s/Hz) based on HD scheme satisfy
@@ III Achievable Spectral Efficiency With Adaptive Optical Beam Transmission
$\mathrm{S}_{\scriptscriptstyle\mathrm{HD}}(\lambda)\,$ $=\,\,\ln\left(1+\dfrac{\lambda P(\lambda)}{N_{\mathrm{hd}}}\right)$ (12)
@@
where $\lambda:=I$, $P(\lambda):=|x_{\mathrm{\scriptscriptstyle HD}}|^{2}$ and $N_{\mathrm{hd}}:=\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}$. The achievable long-term (over blocks) or average spectral efficiency (ASE) is given by
@@
Note that ${P_{\mathrm{avg}}}/{N_{\mathrm{hd}}}$ represents the average transmit power normalized by the average noise power, which will henceforth be termed the transmit SNR and denoted by $\mathrm{SNR}$. The optimal beam power adaptation $P^{\,*}$ maximizing (13) is
$\dfrac{P^{\,*}(\lambda)}{N_{\mathrm{hd}}}\,=\,\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}$ (14)
with $\mu$ chosen so that the optical power constraint
$\mathbb{E}\biggl[\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}\biggr]\,=\,\mathrm{SNR}$ (15)
@@
[…]
@@
In evaluating the integrals $I_{1},\,I_{2}$ and $I_{3}$, we will frequently apply an integral identity involving the Meijer-G function [36, Eq. (9.301)], which is reproduced below (see [47] for additional details).
@@
$G^{m,\,n+1}_{\!p+1,\,q+1}\hskip-2.5pt\left(x\middle|\hskip-3.50006pt\begin{array}[]{c}1,j\!+a_{1},\ldots,j\!+a_{n},j\!+a_{n+1},\ldots,j\!+a_{p}\\ j\!+b_{1},\ldots,j\!+b_{m},\!0,j\!+b_{m+1},\ldots,j\!+b_{q}\end{array}\hskip-5.0pt\right)\!.$ (23)
@@ III-A Exact ASE With and Without Pointing Errors
For the coherent HD optical channel subjected to GG turbulence with pointing errors as described in (9), and with perfect channel state information (CSI) at both ends, the exact ASE in (13) is simplified to
@@
where $\psi(\cdot)$ is the Euler’s digamma function, and $\mu$ is determined from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}$.
@@
$\phantom{xxxxxxxxxxxx}\mathbb{E}\left[\lambda^{t}\ln(\lambda)\right]\bigl|_{\,t=0}\,\,=\,\frac{d}{dt}\left(\mathbb{E}\left[\lambda^{t}\right]\right)\biggl|_{\,t=0}$
$\text{and hence}\phantom{xxxxxx}I_{1}\,=\,\frac{d}{dt}\left(\int_{0}^{\infty}\!\!\lambda^{t}f_{\lambda}(\lambda)d\lambda\right)\biggl|_{\,t=0}\,.$ (29)
@@
$\phantom{xx}\lim_{x\to 0}\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=0,\,\,\,\text{and}$ (35)
$\phantom{xx}\lim_{x\to\infty}\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=\frac{\Gamma\!\left(a\!+t\right)\!\Gamma\!\left(b\!+t\right)}{\xi^{2}\!+t},$ (38)
@@
We apply the integration-by-parts technique treating $\ln(\cdot)$ as the first function and the Meijer-G function as the second to get
@@
$\,\,\,\,\,\,\,\,\,-G_{3,5}^{3,2}\left(\dfrac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\biggr]\,\biggl|_{\,\lambda=0}^{\,\lambda=\mu}\,\,\times\,\,\,\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}.$ (49)
@@
$\,\,\,\,\,\,\,\,\,-\left.G_{3,5}^{3,2}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\right]\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}\cdot$ (61)
@@
= $=\,1-\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right).$ (68)
@@
The ASE with dynamic beam power adaptation over the GG turbulence channel (but without pointing error) is a useful baseline for comparison as follows.
@@
For the coherent HD optical channel subjected to GG turbulence ‘without pointing error’ as described in (2), and with perfect CSI at both ends, the exact ASE is given by
@@
where $\mu$ is solved from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\mathrm{SNR}$.
@@
In the limiting case where $A_{\mathrm{\scriptscriptstyle 0}}\to 1$ and $\xi^{2}\to\infty$, the fading distribution converges to the pure $G$ distribution. Substituting these limits into (1) and applying the identity from [36, Eq. (9.31.1)] completes the proof of the theorem. ∎
To the best of the authors’ knowledge, the proposed exact ASE solutions, Theorems 1 and Corollary 1, for the GG turbulence channels with and without pointing error are new.

### v10

@@ II-B Noise Sources
Photodetection in FSO receivers is fundamentally limited by two noise mechanisms: i) shot noise associated with the quantum nature of photoelectric detection [38], and ii) thermal noise originating from the receiver electronics. Shot noise arises due to both the information-bearing optical field and ambient background radiation collected by the receiver aperture, with a variance proportional to the total incident optical power. Importantly, both ambient shot noise and thermal noise are signal-independent, as they remain statistically uncorrelated with the information-carrying optical signal [39].
In coherent FSO receivers, the received optical field is mixed with a strong optical local oscillator (OLO) signal prior to photodetection. This mixing process introduces an OLO-driven shot-noise component in addition to signal-dependent shot noise, background-induced shot noise, and thermal noise. By appropriately selecting a sufficiently large OLO power, the OLO-induced shot noise can be made to dominate all other noise contributions. Operation in this shot-noise-limited regime is widely regarded as the canonical and practically relevant mode for ideal coherent receivers [7][39][40], and forms the basis for most performance and capacity analyses. By contrast, a direct-detection (DD) receiver produces a photocurrent proportional to the received optical intensity and is subject to signal-dependent shot noise, background-induced shot noise, and thermal noise, with the dominant impairment determined by their relative strengths.
@@
Remark on the Statistical Characterization of Shot Noise: Under low received photon counts—as is typical in inter-satellite and ground-to-satellite optical links—the shot noise is accurately modeled using a Poisson distribution [41]. In contrast, for high photon-count regimes, as encountered in most terrestrial FSO applications, a Gaussian approximation is sufficiently accurate. For a rigorous justification of the Gaussian model for high-intensity shot noise, refer [41][42].
II-C Coherent Heterodyne FSO Receiver Model
@@
Let $x_{\mathrm{\scriptscriptstyle HD}}\in\mathbb{C}$ denote the transmitted symbol for the coherent FSO system respectively. We remind the reader that the baseband (information) signal in a coherent modulation scheme represents a complex optical field, whereas the baseband signal in an intensity modulation scheme represents optical intensity. The transmitted optical power is constrained by a fixed long-term constraint denoted by $\mathbb{E}\left[|x_{\mathrm{\scriptscriptstyle HD}}|^{2}\right]\leqslant P_{\mathrm{avg}}$. As discussed in Section I, owing to millisecond-scale coherence times and coherence bandwidths on the order of hundreds of gigahertz, a frequency-flat block-fading model is appropriate for terrestrial FSO channels subject to time-varying irradiance fluctuations: each fading state is assumed to span an identical duration corresponding to a fixed coherence time, with i.i.d. fading realizations across blocks. The block-wise channel state is assumed to be quasi-ideally known at both the transmitter and the receiver. In practice, the channel state is estimated at the receiver once at the beginning of each coherence block and conveyed to the transmitter via a feedback link. While channel estimation, processing, and CSI feedback inevitably incur non-zero latency, the quasi-ideal CSI assumption adopted here requires that the overall feedback delay be sufficiently small—on the order of sub-millisecond latency—so as to remain negligible relative to the channel coherence time, thereby enabling timely block-wise signal adaptation at the transmitter. Adhering to this understanding, we develop the received signal model for the coherent FSO communication system employing HD detection. Assuming ideal synchronous HD detection, the received complex baseband symbol is described by
$y_{\mathrm{\scriptscriptstyle HD}}\,=\,h\,x_{\mathrm{\scriptscriptstyle HD}}\,+\,\,w_{\mathrm{\scriptscriptstyle OLO}}\,$, (10)
@@
where $y_{\mathrm{\scriptscriptstyle HD}}\ \in\ \mathbb{C}$ represents its output, $h\in\mathbb{C}$ is the fluctuation in the received optical field and $w_{\mathrm{\scriptscriptstyle OLO}}\sim\mathcal{CN}(0,\sigma^{2}_{\mathrm{\scriptscriptstyle OLO}})$ is the OLO-induced shot noise. As noted earlier, the OLO shot noise is the only dominant noise of consequence, and the total AWGN from all sources other than the OLO is negligible [43, Ch. 3]. For the baseband model in (10), the received electrical SNR conditioned on $h$ is
$\frac{IP(I)}{\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}}\,\cdot$ (11)
Here, $I\triangleq|h|^{2}$ denotes the channel (intensity) gain, and $P(I)\triangleq\mathbb{E}\!\left[|x_{\mathrm{\scriptscriptstyle HD}}|^{2}\,\big|\,I\right]$ represents the transmit optical power allocated according to the channel state $I$.
Since the average transmit power budget and the OLO shot-noise variance are fixed system parameters, it is convenient to express power in normalized form by scaling $P_{\mathrm{avg}}$ with $\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}$, a ratio that naturally arises in the subsequent derivation.
Definition 1.
The normalized transmit power is defined as
$\rho\triangleq\frac{P_{\mathrm{avg}}}{\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}}\,\cdot$ (12)
Under ideal coherent FSO conditions—including unit channel gain (i.e., no turbulence or pointing errors and full beam capture) and ideal optical-to-electrical conversion—the entire transmitted optical power is recovered at the receiver, and $\rho$ coincides with the received electrical SNR.
Remark 2.
The adopted coherent FSO framework for (10) assumes that the OLO-driven receiver shot noise is the dominant impairment. Although transmitter-induced noise sources (e.g., laser RIN, modulation noise, and optical amplifier ASE noise) originate at the transmitter, their effective contribution must be assessed at the receiver after propagation, optical mixing and detection. The present analysis assumes that these contributions are small compared to the OLO shot noise at the receiver, and are therefore not explicitly modeled. Accordingly, the parameter $\rho$ represents a nominal transmit-power normalization referenced to the OLO shot-noise variance at the receiver, and should not be interpreted as a hardware-constrained transmit SNR.
Any performance variations arise exclusively from the channel statistics and their exploitation via channel-dependent power adaptation observing the budget $\mathbb{E}[P(I)]=P_{\mathrm{avg}}$. Accordingly, $\rho$ is adopted throughout this paper as the channel-independent system resource parameter for ASE evaluation and comparison across different turbulence and pointing-error regimes.
@@
III Achievable Spectral Efficiency With Optimal Adaptive Coherent FSO Transmission
In IM-DD FSO systems, information is mapped onto the intensity of the transmitted optical field, which is inherently non-negative. In constrast, the baseband symbol $x_{\mathrm{\scriptscriptstyle HD}}$ in (10) utilizes two degrees of freedom. Since $w_{\mathrm{\scriptscriptstyle OLO}}$ is AWGN, the channel conditioned on the fading state admits the classical Shannon capacity expression. Accordingly, for the instantaneous received SNR in (11), the achievable spectral efficiency (in nats/s/Hz) of the coherent FSO channel in (10) is
@@ III Achievable Spectral Efficiency With Optimal Adaptive Coherent FSO Transmission
$\mathrm{S}_{\scriptscriptstyle\mathrm{HD}}(\lambda)\,$ $=\,\,\ln\left(1+\dfrac{\lambda P(\lambda)}{\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}}\right)\,\mathpunct{\raisebox{2.15277pt}{,}}$ (13)
@@
where $\lambda\triangleq I$ denotes the block fading state. The maximum long-term average (over blocks) or simply average spectral efficiency (ASE) is defined as
@@
The optimal beam power adaptation $P^{\,*}$ maximizing (14) is
$\dfrac{P^{\,*}(\lambda)}{\sigma_{\mathrm{\scriptscriptstyle OLO}}^{2}}\,=\,\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}\,\mathpunct{\raisebox{2.15277pt}{,}}$ (15)
with $\mu$ chosen so that the (normalized) optical power constraint
$\mathbb{E}\biggl[\left(\dfrac{1}{\mu}-\dfrac{1}{\lambda}\right)^{+}\biggr]\,=\,\rho$ (16)
@@
The integral in (17) can be divided into three parts as follows:
@@
In evaluating the $I_{1},\,I_{2}$ and $I_{3}$ integrals, we will frequently apply an integral identity involving the Meijer-G function [32, Eq. (9.301)], which is reproduced below (see [44] for additional details).
@@
$G^{m,\,n+1}_{\!p+1,q+1}\hskip-2.5pt\left(x\middle|\hskip-3.99994pt\begin{array}[]{c}1,j\!+a_{1},\ldots,j\!+a_{n},j\!+a_{n+1},\ldots,j\!+a_{p}\\ j\!+b_{1},\ldots,j\!+\!b_{m},\!0,j\!+b_{m+1},\ldots,j\!+b_{q}\end{array}\hskip-5.0pt\right)\!\cdot$ (24)
@@ III-A Exact ASE With and Without Pointing Errors
For the adaptive coherent terrestrial FSO channel in (10), with intensity gain $I=|h|^{2}$ modeled by a composite fading distribution accounting for GG turbulence and pointing errors as characterized in (9), and assuming perfect channel state information (CSI) at both the transmitter and the receiver, the exact ASE in (14) simplifies to
@@
where $\psi(\cdot)$ is the Euler’s digamma function, and $\mu$ is determined from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\rho$.
@@
$\phantom{xxxxxxxxxxxx}\mathbb{E}\left[\lambda^{t}\ln(\lambda)\right]\bigl|_{\,t=0}\,\,=\,\frac{d}{dt}\left(\mathbb{E}\left[\lambda^{t}\right]\right)\biggl|_{\,t=0}\,\,$,
$\text{and hence}\phantom{xxxxx}I_{1}\,=\,\frac{d}{dt}\left(\int_{0}^{\infty}\!\!\lambda^{t}f_{\lambda}(\lambda)d\lambda\right)\biggl|_{\,t=0}\,\,.$ (30)
@@
$\phantom{x}\,\,\lim_{x\to 0}\,\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=0,\,\,\,\text{and}$ (36)
$\phantom{x}\,\,\lim_{x\to\infty}\,G_{2,4}^{3,1}\!\left(x\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\xi^{2}\!+t+1\\ \!\xi^{2}\!+t,\!a\!+t,\!b\!+t,\!0\\ \end{array}\hskip-5.0pt\right)=\frac{\Gamma\!\left(a\!+t\right)\!\Gamma\!\left(b\!+t\right)}{\xi^{2}\!+t}\,\mathpunct{\raisebox{2.15277pt}{,}}$ (39)
@@
Applying integration by parts with $\ln(\cdot)$ as the first function and the Meijer-G function as the second yields
@@
$\,\,\,\,\,\,\,\,\,-G_{3,5}^{3,2}\left(\dfrac{ab\lambda}{A_{\mathrm{\scriptscriptstyle 0}}}\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\biggr]\,\biggl|_{\,\lambda=0}^{\,\lambda=\mu}\,\,\times\,\,\,\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,\cdot$ (50)
@@
$\,\,\,\,\,\,\,\,\,-\left.G_{3,5}^{3,2}\left(\dfrac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,1,\xi^{2}+1\\[1.00006pt] \xi^{2},a,b,0,0\end{array}\hskip-3.50006pt\right)\right]\,\,\times\,\,\,\dfrac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,\cdot\,\,\,\,$ (62)
@@
= $=\,1-\frac{\xi^{2}}{\Gamma(a)\Gamma(b)}\,G_{2,4}^{3,1}\left(\frac{ab}{A_{\mathrm{\scriptscriptstyle 0}}}\mu\,\biggr\rvert\hskip-2.5pt\begin{array}[]{c}1,\,\xi^{2}+1\\[1.00006pt] \xi^{2},\,a,\,b,\,0\\ \end{array}\hskip-3.50006pt\right)\,\cdot$ (69)
@@
The exact ASE performance with optimal beam power adaptation over the GG turbulence channel, excluding pointing errors, serves as a useful baseline for comparison.
@@
For the adaptive coherent terrestrial FSO channel in (10), with intensity gain $I=|h|^{2}$ modeled by a GG distribution without pointing errors as characterized in (2), and assuming perfect channel state information (CSI) at both the transmitter and receivers, the exact ASE is given by
@@
where $\mu$ is solved from $\int_{\mu}^{\infty}(1/\mu-1/\lambda)f_{{\lambda}}(\lambda)d\lambda\,=\,\rho$.
@@
In the limiting case where $A_{\mathrm{\scriptscriptstyle 0}}\!\to\!1$ and $\xi^{2}\!\to\!\infty$, the composite fading distribution in (9) converges to the GG distribution in (2). Substituting these into (1) and applying the identity from [32, Eq. (9.31.1)] completes the proof. ∎
To our knowledge, Theorem 1 and Corollary 1 provide novel exact ASE expressions for adaptive coherent terrestrial FSO links under GG turbulence, with and without pointing errors.